OpenStudy (anonymous):

Solve for the variable in the following proportion. 3.6/y = 1.2/2

OpenStudy (anonymous):

first solve 1.2/2=0.6. Then multiply both sides by y: \[3.6/y(y)=0.6(y)\] the y's cancel out on the side with 3.6 and now you have: \[3.6=0.6y\] \[y=6\]

OpenStudy (anonymous):

thats helpful thanks

OpenStudy (anonymous):

Your welcome